OpenStudy (anonymous):

When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week?

OpenStudy (anonymous):

@magepker728

OpenStudy (anonymous):

(40+x)=10+(4x) -x = 40=10+3x 30=3x x=10 verify 10+40=50 10+(4x10)=50 x=10

OpenStudy (anonymous):

Thank you so much!